OpenStudy (anonymous):

I X-3 I=5X+13 solve the equation....

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

-x+3 = 5x + 13 6x = -10 x = -5/3 or x-3 = 5x + 13 4x = -16 x = -4 there are two roots because the | | ie the absolute value of whats inside these lines. this can be either negative or positive

OpenStudy (radar):

\[|x-3|=5x+13\]\[x-3=\pm(5x+13)\] x-3=5x+13 x-5x=13+3 -4x=16 x=-4 One value x-3=-5x-13 x+5x=-13+3 6x=-10 x=-10/6=-5/3 other value

OpenStudy (anonymous):

hi radar how are you?

OpenStudy (anonymous):

kaitlyn - thats a celtic name

OpenStudy (radar):

Doing fine for an old man jimmy, hope u r 2.

OpenStudy (anonymous):

yes - ok thanx